Beta airlines awards frequent flyer miles that are equal to the number of miles traveled, rounded to the nearest hundred. If an airplanes average speed for 5.5 hour nonstop flight from Tampa to Los Angeles is 470 MPH, how many frequent flyer miles are earned from the flight?
Answer by jim_thompson5910(35256)


Speed = 470 mph
Time = 5.5 hours

Distance = Speed*Time
Distance = 470*5.5
Distance = 2585 miles

The one way distance from Tampa to Los Angeles is 2585 miles. Round this to the nearest hundred and we get 2600. Put another way, the value 2585 is closer to 2600 than it is to 2500.

Compare the differences
2600-2585 = 15
2585-2500 = 85
And note the smaller difference when we subtract 2600 and 2585

Therefore, 2600 frequent flyer miles are earned on this flight.